OpenStudy (anonymous):

Evaluate:

OpenStudy (anonymous):

\[\log _{4}16\]

OpenStudy (anonymous):

Log base 4 16 right?

OpenStudy (anonymous):

Yupp

OpenStudy (anonymous):

k

OpenStudy (anonymous):

16 = 4^x 4^2 = 4^x x = 2

OpenStudy (anonymous):

Got it..thanks

OpenStudy (anonymous):

k good luck